# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ C.3 Cumulative Prospect Theoretic RL
Motivated by Prospect Theory and its sibling CPT [Kahneman and Tversky, 1979, Tversky and Kahneman, 1992, Barberis, 2013], L.A. et al. [2016] first proposed to combine CPT with RL to obtain a better model for human decision making. Following this first research effort, only few isolated works [Borkar and Chandak, 2021, Ramasubramanian et al., 2021, Ethayarajh et al., 2024] considered a similar CPT-RL setting. In particular, Borkar and Chandak [2021] proposed and analyzed a Q-learning algorithm for CPT policy optimization. Ramasubramanian et al. [2021] further developed value-based algorithms for CPT-RL by estimating the CPT value of an action in a given state via dynamic programming. More precisely, they were concerned with maximizing a sum of CPT value period costs which is amenable to dynamic programming. In contrast to their accumulated CPT-based cost (see their remark 1), our CPT policy optimization problem formulation is different: we maximize the CPT value of the return of a policy (see (CPT-PO)). In particular, this objective does not enjoy an additive structure and hence does not satisfy a Bellman equation. Moreover, their work relying on value-based methods is restricted to finite discrete state action spaces. Our PG algorithm is also suitable for continuous state action settings as we demonstrate in our experiments. More recently, Ethayarajh et al. [2024] incorporated CPT (without probability distortion) into RL from human feedback for fine-tuning large language models. CPT has also been recently exploited for multi-agent RL [Danis et al., 2023]. Our work is complementary to this line of research, especially to L.A. et al. [2016] and its extended version Jie et al. [2018] which are the most closely related work to ours. While their algorithm design makes use of simultaneous perturbation stochastic approximation (SPSA) [Spall, 1992] using only zeroth order information, we rather propose a PG algorithm exploiting first-order information thanks to our special problem structure involving the CPT value of a cumulative sum of rewards. See section 4 for further details regarding this comparison.
@@
We refer the reader to Appendix F.1 for a summarizing diagram illustrating the relationships between CPT-RL, convex RL and risk-sensitive RL.
Appendix D Applications of CPT: From Prior Work in Stateless Settings to CPT-RL
@@
    ‘We conduct a standardized survey on risk preferences in 53 countries worldwide and estimate cumulative prospect theory parameters from the data. The parameter estimates show that significant differences on the cross-country level are to some extent robust and related to economic and cultural differences. In particular, a closer look on probability weighting underlines gender differences, economic effects, and cultural impact on probability weighting.’
D.1 Energy: Renovation and Home Energy Management
@@
    ‘CPT correctly predicts the decisions of 86 % of homeowners to renovate their homes to be energy efficient or not. EUT, on the other hand, overestimates the number of decisions to renovate: it incorrectly predicts retrofit for 52 % of homeowners who did not renovate for energy efficiency reasons. Using the estimated parameters of CPT, the cognitive biases of reference dependence, loss aversion, diminishing sensitivity, and probability weighting can be clearly seen for different target groups.’
@@
    ‘Smart home is a small but an important energy segment that has a significant potential to implement authentic energy policies, where human is a major decision-maker in the home energy management dilemma. Therefore, humans’ emotions and tendencies plays a vital role as the End-User’s daily decisions. In this paper, we develop a behavioral home energy management model based on time-driven prospect theory incorporating energy storage devices, distributed energy resources, and smart flexible home appliances. […] The results of the simulation studies show that the End-User’s satisfaction in the proposed home energy management behavioral structure will be increased substantially compared with the conventional monetary home energy management models.’
D.2 Security: Building Evacuation
@@
    ‘Understanding the performance of decision-making behavior in building evacuation is essential for predicting pedestrian dynamics, designing appropriate facility safety management, optimizing emergency management strategies, and reducing the impact of disasters. While many pedestrian movement models have been developed based on the hypothesis of rational and strategic decision-making, only a limited number of works consider individual psychology and irrational behavior. To address this issue, we have successfully integrated the cumulative prospect theory (CPT) into modeling evacuations.’
D.3 Urban Planning and Mobility
@@
    ‘Results show that socio-demographic characteristics, context variables, revenues and psychological concerns are all important factors in explaining parking space owners’ propensity to engage in platform-based shared parking schemes. […] Understanding parking space owners’ propensity to share their parking spaces in relation to their psychological concerns and uncertain conditions is critical to improve shared parking policies. The results of this paper may help designers and planners in the delivery of shared parking services and promote the success and future growth of the shared parking industry.’
@@ D.4 Finance
    ‘This paper aims to provide empirical evidence for using the prospect theory (PT) basic assumptions in the Argentine context. Mainly, this study analysed the financial decision-making process in students of the economic-administrative academic area of two universities, one public and one private, in Córdoba. […] The empirical results provided evidence on the effects of certainty, reflection and isolation in both universities, concluding that the study participants make financial decisions in situations of uncertainty based more on PT than on expected utility theory. This study contributes to the empirical evidence in a different Latin-American context, confirming that individuals make financial decisions based on the PT […].’
D.5 CPT-RL applications
@@
  • Electricity management. Please see simulations in the main part (section 5 and appendix I.6) in a simple example setting to illustrate our methodology.
  • Finance: portfolio optimization, risk management, behavioral asset pricing (e.g. influence of investor sentiment on price dynamics via e.g. over-weighting of low-probability events, including their preferences). For recent applications of CPT to finance, we refer the reader to a recent paper Luxenberg et al. [2024] using CPT for portfolio optimization (in a stateless static setting). We also applied our methodology to financial trading (see Appendix I.7).
@@
Many other meaningful human-centric applications are yet to be explored, including: such as
  • Legal and ethical decision making,
  • Cybersecurity,
  • Human-robot interaction,
Appendix E CPT-RL and Trajectory-Based Reward RL as Preference Learning Paradigms
@@
Appendix F More about CPT Values and CPT Policy Optimization
F.1 Positioning CPT-RL in the literature
@@
F.2 CPT value examples
Setting Utility function $w^{+}$ $w^{-}$
@@ F.2 CPT value examples
CVaR* (Balbás et al. [2009]) Identity function $1-w^{-}(1-t)$$\begin{cases}\frac{x}{1-\alpha}&\text{if }0\leq x<1-\alpha,\\ 1&\text{if }1-\alpha\leq x\leq 1\end{cases}$
VaR* [Balbás et al., 2009] Identity function $1-w^{-}(1-t)$$\begin{cases}0&\text{if }0\leq x<1-\alpha,\\ 1&\text{if }1-\alpha\leq x\leq 1\end{cases}$
@@
F.3 Proof: CVar, Var and distortion risk measures are CPT values
@@
[…]
Since $g(\mathbb{P}(-X>x))=g(\mathbb{P}(-X\geq x))$ almost everywhere (in a measure theoretic sense) on [0,+([0,$\infty$([ 0 , + ∞ (, and $g$ is bounded, we obtain:
[…]
@@
F.4 Simple examples and further insights about CPT
@@
Let us illustrate this in a simple example borrowed from Ramasubramanian et al. [2021] (example 2 in section IV therein) for the purpose of our exposition. Consider a game where one can either earn \$$\$100$ with probability (w.p.) 1 or earn $10000$ w.p. $0.01$ and nothing otherwise. A human might rather lean towards the first option which gives a certain gain. In contrast, if the situation is flipped, i.e., a loss of $100$ w.p. 1 versus a loss of \$$\$10000$ w.p. 0.01, then humans might rather choose the latter option. In both settings, the expected gain or loss has the same value ($100$). The CPT paradigm allows to model the tendency of humans to perceive gains and losses differently. Moreover, the humans tend to deflate high probabilities and inflate low probabilities [Tversky and Kahneman, 1992, Barberis, 2013]. For instance, as exposed in L.A. et al. [2016], humans might rather choose a large reward, say $1$ million dollars w.p. $10^{-6}$ over a reward of $1$ w.p. $1$ and the opposite when rewards are replaced by losses.
F.5 Connection to General Utility RL and Convex RL in finite trials
In this section, we elaborate in more details on one of the connections we noticed (and mentioned in related works) between our (CPT-PO) problem of interest and the literature of generality utility RL.
@@ F.5 Connection to General Utility RL and Convex RL in finite trials
Setting the probability distortion function $w$ to be the identity, our (CPT-PO) problem becomes (EUT-PO), i.e.:
@@
F.6 Choice of the utility function
F.7 Monotonicity assumption on the utility function
We provide a few comments about the monotonicity assumption we make in Theorem 4 for instance, and its implications for the remaining results of the paper. We argue below that this is not a restriction and it is not fundamentally needed for our algorithm:
  • We focus on the monotone setting because typical human behavior tends to prefer better outcomes over worst ones, e.g. higher gains over smaller ones and smaller losses over larger losses. This is captured by the monotonicity assumption on the utility function: utility increases with increasing gains and decreases with increasing losses. This is a fundamental assumption in economics and decision theory which is consistent with how humans evaluate outcomes. Nevertheless, there are certainly cases where this assumption might not hold.
  • Strict monotonicity is needed in some of the proofs of our results, mainly Theorem 4, see e.g. proof of (1) implies (2) and proof of (5) implies (2). However, it is not formally required to derive our policy gradient theorem. Note also that even Proposition 7 for estimating the integrals does not require it as soon as denote the right quantiles as we defined them in the proposition. See point 3 below for further comments.
  • Mere monotonicity is mainly useful in our algorithm to guarantee that the utility values are sorted in the same order as the returns obtained from the sampled trajectories (see quantile estimation computation step of the algorithm). This leads to a simple computation of the quantiles of the utilities: Once the returns are sorted, the utilities are also sorted in the same way and quantiles can be read from the sorted list. If monotonicity does not hold anymore, one has to be more careful about this computation and adapt the step of the algorithm accordingly by simply computing the quantiles of the utilities using the right sorting of these (which would be different from the sorting of the returns). In that case, sorting the returns is not needed and one only needs to sort the utilities (utility function applied to the returns). Apart from this technical detail, we do not foresee any major impediment to using the algorithm without the strict monotonicity assumption.
Appendix G Proofs for Section 3
G.1 Unwinding MDPs for CPT-RL
In this section, we describe an equivalent MDP construction that will be used in some of our proofs such as for Proposition 3. For any CPT-MDP $(\mathcal{S},\mathcal{A},r,P)$ with utility function $\mathcal{U}$, we can formally define an equivalent ‘unwinded’ MDP [^5] This terminology is not standard, we adopt it here to describe our approach. that can be solved using classical RL techniques. For any state $s\in\mathcal{S}$ is the original MDP and any timestep $t\leq H-1$ with cumulative reward $\sum_{k=0}^{t}r_{k}$, we associate a state $\tilde{s}:=(s,t,\sum_{k=0}^{t}r_{k})$ and the rewards in the unwinded MDP are adjusted as to reflect the difference in utility between two consecutive states:
$\tilde{r}_{t}=\mathcal{U}\left(\sum_{k=1}^{t+1}r_{k}\right)-\mathcal{U}\left(\sum_{k=1}^{t}r_{k}\right)\,.$ (8)
We observe that all the information needed at any given timestep to take a decision on the next action to take is contained in $\tilde{s}$. This implies that any CPT-value that can be achieved by a non-Markovian strategy on the original MDP can also be achieved by a Markovian policy on the unwinded MDP.
The reader might notice that the size of the unwinded MDP grows with the horizon length and might blow up depending on the original MDP structure. As a consequence, learning in this unwinded MDP might become intractable. If the original MDM can be represented as a finite directed acyclic graph, the unwinded MDP is also a finite directed acyclic graph. If the underlying MDP contains a cycle, even if it is finite, its unwinded version may contain an infinite number of states. In the case of a stochastic tree MDP, the unwinded MDP has the exact same shape as the original one.
Note that we will only be using the unwinded MDP as a theoretical construction to prove some of our results and we do not perform any learning task in this unwinded MDP.
G.2 Proof of Proposition 2
To prove the proposition, we consider a simple MDP with only two states (an initial state and a terminal one) and two actions (A and B). See Fig. 14(a) below. We choose the identity as utility. Action A yields reward 1 with probability 1 and action B yields either 0 or $\frac{3}{2}$ with probability $\frac{1}{2}$ each. We further consider the following probability distortion function $w^{+}:[0,1]\to[0,1]$ defined for every $x\in[0,1]$ as follows:
$w^{+}(x)=\begin{cases}5x&\text{if }x\leq 0.1,\\ \frac{1}{2}+\frac{5}{9}(x-0.1)&\text{otherwise}\,,\end{cases}$ (9)
@@ G.2 Proof of Proposition 2
Figure 9: Problem instance for the proof of Proposition 2.

### v3

@@ D.3 Cumulative Prospect Theoretic RL
Motivated by Prospect Theory and its sibling CPT [Kahneman and Tversky, 1979, Tversky and Kahneman, 1992, Barberis, 2013], L.A. et al. [2016] first proposed to combine CPT with RL to obtain a better model for human decision making. Following this first research effort, only few isolated works [Borkar and Chandak, 2021, Ramasubramanian et al., 2021, Ethayarajh et al., 2024] considered a similar CPT-RL setting. In particular, Borkar and Chandak [2021] proposed and analyzed a Q-learning algorithm for CPT policy optimization. Ramasubramanian et al. [2021] further developed value-based algorithms for CPT-RL by estimating the CPT value of an action in a given state via dynamic programming. More precisely, they were concerned with maximizing a sum of CPT value period costs which is amenable to dynamic programming. In contrast to their accumulated CPT-based cost (see their remark 1), our CPT policy optimization problem formulation is different: we maximize the CPT value of the return of a policy (see CPT-PO). In particular, this objective does not enjoy an additive structure and hence does not satisfy a Bellman equation. Moreover, their work relying on value-based methods is restricted to finite discrete state action spaces. Our PG algorithm is also suitable for continuous state action settings as we demonstrate in our experiments. More recently, Ethayarajh et al. [2024] incorporated CPT (without probability distortion) into RL from human feedback for fine-tuning large language models. CPT has also been recently exploited for multi-agent RL [Danis et al., 2023]. Our work is complementary to this line of research, especially to L.A. et al. [2016] and its extended version Jie et al. [2018] which are the most closely related work to ours. While their algorithm design makes use of simultaneous perturbation stochastic approximation (SPSA) [Spall, 1992] using only zeroth order information, we rather propose a PG algorithm exploiting first-order information thanks to our special problem structure involving the CPT value of a cumulative sum of rewards. See section 3 for further details regarding this comparison.
@@
We refer the reader to App. G.2 for a summarizing diagram illustrating the relationships between CPT-RL, convex RL and risk-sensitive RL.
Appendix E Applications of CPT: From Prior Work in Stateless Settings to CPT-RL
@@
E.1 Energy: Renovation and Home Energy Management
@@
E.2 Security: Building Evacuation
@@
E.3 Urban Planning and Mobility
@@ E.4 Finance
E.5 Example: Personalized Treatment for Pain Management
We illustrate our CPT-RL problem formulation with a concrete example in healthcare to give the reader more intuition about the different features of CPT-RL, its importance in applications when human perception and behavior matter and its differences compared to risk-sensitive RL. The goal is to help a physician manage a patient’s chronic pain by suggesting a personalized treatment plan over time. The challenge here is to balance pain relief and the risk of opioid dependency or other side effects that might be due to the treatment, i.e. short-term relief and longer term risks. The idea is to train a CPT-RL agent to help the physician.
1. Why RL? (a) The physician needs to adjust treatment at each time step depending on the patient’s reported pain level as well as the observed side effects. This is relevant to dynamic treatment regimes in general (such as for chronic diseases, see e.g. Yu et al. [2021] for a survey) in which considering delayed effects of treatments is also important and RL does account for such effects. (b) Decisions clearly impact the patient’s immediate pain relief, dependency risks in the future and their overall health condition. A state is described by e.g. current pain level, dependency risk and side effect severity. Actions are treatments, e.g. no treatment, opioid or alternative treatment.
2. Why CPT-RL? Patients and clinicians make decisions influenced by psychological biases. We illustrate the importance of each one of the three features of CPT in section 2 (reference point, utility and probability distortion weight functions) via this example:
  • Utility transformation: Patients might often show a loss averse behavior, i.e., they might perceive pain increase or withdrawal symptoms as worse than equivalent gains in pain relief. Note here that loss aversion should not be confused with risk aversion [Schmidt and Zank, 2005]. In short, loss aversion can be defined as a cognitive bias in which the emotional impact of a loss is more intense than the satisfaction derived from an equivalent gain. For instance, in our example, a 2-point increase in pain might be seen as much worse than a 2-point reduction even if the change is the same in absolute value. This loss aversion concept is a cornerstone of Kahneman and Tversky’s theory. In contrast, risk aversion rather refers to the rational behavior of undervaluing an uncertain outcome compared to its expected value. Risk sensitive approaches might be less adaptive to a patient’s subjective preferences if they deviate from objective risk assessments.
3. CPT-RL vs Risk-averse RL. In terms of policies, risk-averse RL would favor non-opioid treatments unless extreme pain levels make opioids justifiable. In contrast, CPT-RL policies would prescribe opioids if pain significantly exceeds the patient’s reference point. As dependency risk increases, CPT-RL policies would transition to non-opioid treatments as a consequence of overweighting the probability of rare catastrophic outcomes. Notably, CPT-RL policies can oscillate between risk-seeking (to address high pain) and risk-averse (to avoid severe side effects). In contrast, a risk-sensitive agent focuses on minimizing variability in health states and dependency risks and would likely avoid opioids in most cases unless pain levels become extreme. Such risk-sensitive policies favor stable strategies (e.g., consistent non-opioid use), prioritizing low variance in patient outcomes.
E.6 Further Applications of CPT-RL
@@
  • Electricity management. Please see simulations in the main part (section 5 and App. K.7) in a simple example setting to illustrate our methodology.
  • Finance: portfolio optimization, risk management, behavioral asset pricing (e.g. influence of investor sentiment on price dynamics via e.g. over-weighting of low-probability events, including their preferences). For recent applications of CPT to finance, we refer the reader to a recent paper Luxenberg et al. [2024] using CPT for portfolio optimization (in a stateless static setting). We also applied our methodology to financial trading (see App. K.8).
@@
Many other meaningful human-centric applications are yet to be explored, including legal and ethical decision making, cybersecurity and human-robot interaction.
Appendix F CPT-RL and Trajectory-Based Reward RL as Preference Learning Paradigms
@@
Appendix G More About CPT Values and CPT Policy Optimization
G.1 More about Optimal Policies in CPT-RL
The need for stochastic policies. We start our discussion by pointing out a stark difference between optimal policies in standard MDPs and CPT-PO. While there exists an optimal deterministic stationary policy for MDPs (see e.g. Thm. 6.2.10 in Puterman [2014]), this is not the case in general for CPT-PO. Indeed, there does not always exist an optimal policy for CPT-PO in $\Pi_{NM}^{D}\,.$ In general, stochasticity of the policy is essential in solving our CPT-RL problem. To see this, consider an example built around a function $w^{+}$ that puts special emphasis on the $10\%$ of the best outcomes. In this case, the optimal policy needs to be randomized to take advantage of this and obtain the highest returns with some probability without suffering from bad outcomes by deterministically committing to this riskier strategy (see App. H.1 for a proof).
Importance of probability weighting. When setting the probability weight distortion function $w$ to the identity, i.e. when considering the particular case EUT-PO of CPT-PO, it appears that an optimal policy is not necessarily stochastic. Indeed, it has been shown that there exists an optimal policy for EUT-PO in $\Pi_{\Sigma,NS}^{D}$ (see e.g. Theorem 1 in Bäuerle and Rieder [2014]). Therefore, the need for stochasticity in the optimal policy is clearly due to the probability distortions in the CPT value. The aforementioned result allows to safely restrict the policy search to $\Pi_{\Sigma,NS}$ which is a much smaller policy space than the set of non-Markovian policies $\Pi_{NM}$. The fact that an optimal deterministic policy exists is a fundamental difference with the general CPT-PO setting. Whether there are specific weight functions (apart from the identity) for which there always exist a deterministic optimal policy remains an open question left for future work.
The need for non-Markovian policies. We now ask the next natural question: Can we further restrict our policy search to a smaller policy class compared to $\Pi_{\Sigma,NS}\,$? In particular, are there specific utility functions for which the resulting EUT-PO problem has optimal Markovian policies? Bäuerle and Rieder [2014] provide a positive answer by establishing a precise characterization of such utility functions which turn out to be either affine or exponential (when $\mathcal{U}$ is continuous and increasing). This highlights the role of the (nonlinear) utility functions on the nature of optimal policies. However, these results cannot be extended to CPT-PO in general.
G.2 Positioning CPT-RL in the literature
@@
G.3 CPT value examples
Setting Utility function $^{+} 𝒘w^{-}$
@@ G.3 CPT value examples
[…]
@@
G.4 Proof: CVaR, Var and distortion risk measures are CPT values
@@
$\rho_{g}(X)$ $=\int_{-\infty}^{0}\tilde{g}(F_{-X}(x))dx-\int_{0}^{+\infty}g(1-F_{-X}(x))dx$
$=-\int_{+\infty}^{0}\tilde{g}(F_{-X}(-x))dx-\int_{0}^{+\infty}g(1-F_{-X}(x))dx$
$=\int_{0}^{+\infty}\tilde{g}(F_{-X}(-x))dx-\int_{0}^{+\infty}g(1-F_{-X}(x))dx$
$=\int_{0}^{+\infty}\tilde{g}(\mathbb{P}(-X\leq-x))dx-\int_{0}^{+\infty}g(1-\mathbb{P}(-X\leq x))dx$
$=\int_{0}^{+\infty}\tilde{g}(\mathbb{P}(X\geq x))dx-\int_{0}^{+\infty}g(\mathbb{P}(-X>x))dx\,.$
Since $g(\mathbb{P}(-X>x))=g(\mathbb{P}(-X\geq x))$ almost everywhere (in a measure theoretic sense) on $0,+\infty($, and $g$ is bounded, we obtain:
$\rho_{g}(X)=\int_{0}^{+\infty}\tilde{g}(\mathbb{P}(X\geq x))dx-\int_{0}^{+\infty}g(\mathbb{P}(-X\geq x))dx\,.$
@@
G.5 Simple examples and further insights about CPT
@@
Let us illustrate this in a simple example borrowed from Ramasubramanian et al. [2021] (example 2 in section IV therein) for the purpose of our exposition. Consider a game where one can either earn $\mathdollar 100$ with probability (w.p.) 1 or earn $10000$ w.p. $0.01$ and nothing otherwise. A human might rather lean towards the first option which gives a certain gain. In contrast, if the situation is flipped, i.e., a loss of $100$ w.p. 1 versus a loss of $\mathdollar 10000$ w.p. 0.01, then humans might rather choose the latter option. In both settings, the expected gain or loss has the same value ($100$). The CPT paradigm allows to model the tendency of humans to perceive gains and losses differently. Moreover, the humans tend to deflate high probabilities and inflate low probabilities [Tversky and Kahneman, 1992, Barberis, 2013]. For instance, as exposed in L.A. et al. [2016], humans might rather choose a large reward, say $1$ million dollars w.p. $10^{-6}$ over a reward of $1$ w.p. $1$ and the opposite when rewards are replaced by losses.
G.6 Connection to General Utility RL and Convex RL in finite trials
In this section, we elaborate in more details on one of the connections we noticed (and mentioned in related works) between our CPT-PO problem of interest and the literature of generality utility RL.
@@ G.6 Connection to General Utility RL and Convex RL in finite trials
Setting the probability distortion function $w$ to be the identity, our CPT-PO problem becomes EUT-PO, i.e.:
@@
where $r$ is the reward function seen as a vector in $\mathbb{R}^{|\mathcal{S}|}\,$, $\langle\cdot,\cdot\rangle$ is the standard Euclidean product in $\mathbb{R}^{|\mathcal{S}|}\,.$ Therefore, it appears that the above objective is indeed a functional of the empirical distribution $d_{1}^{\pi}\,.$ Single trial general utility RL is more general than EUT-PO since it does not necessarily consider an additive reward inside the non-linear utility and can accommodate any (convex) functional of the occupancy measure. However, CPT-PO does not appear to be a particular case of single trial convex RL because of the probability distortion function introduced.
G.7 Choice of the utility function
Appendix H Proofs for Section 2.3
H.1 The need for stochastic policies
To prove the result (i.e. the need for stochastic policies), we consider a simple MDP with only two states (an initial state and a terminal one) and two actions (A and B). See Fig. 14(a) below. We choose the identity as utility. Action A yields reward 1 with probability 1 and action B yields either 0 or $\frac{3}{2}$ with probability $\frac{1}{2}$ each. We further consider the following probability distortion function $w^{+}:[0,1]\to[0,1]$ defined for every $x\in[0,1]$ as follows:
@@ H.1 The need for stochastic policies
Figure 9: Problem instance for the proof.